OpenStudy (anonymous):

write in slope intercept and explain y+2=-(x-4)

OpenStudy (anonymous):

Y=X+2 I think. :)

OpenStudy (anonymous):

I'm not too sure how to explain.

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

go on Mathway.com :) it helps!

OpenStudy (anonymous):

y=-x+4-2 y=-x+2 so your slope is -1 y intercept is 2